Flux normalization in GLoBES 2.0.11

J.E Campagne - LAL

June 1, 2005

Introduction

The present situation is the following: we use the files sploptiplusmauro.dat (sploptiminusmauro.dat) which are fluxes (dN(E)) expressed in terms of numbers of neutrino species per 100m^{2} per year and binned every BW F ≡ 0.02GeV and computed at L = 130km.

Internally, GLoBES v2.0.11 uses these fluxes with some a priori and at the end of the day it turns out that:

dN(E) × \frac{1}{L^{2}} × norme2(0) × @norm × ΔT  (1)

with norme2(0) = 5.198927 = 6.0220410^{−12} 295^{2} 9.9203310^{6} the result of a

function in glb_fluxes.c.